(i)           The resultant gravitational force exerted upon the body at this point is zero. Calculate the distance R of the body from the centre of the Earth given that

r = 3.9 × 107 m and ME = 81 MM

1
Expert's answer
2021-03-29T09:04:47-0400

The two forces are in equilibrium, which means the force of attraction between the earth and body (which is r = 3.9 × 107 away from Earth) is the same as the force between the body and mass MM at distance R:


"F_E=F_M,\\\space\\
G\frac{M_Em}{r^2}=G\frac{M_Mm}{R^2},\\\space\\
R=r\sqrt{\frac{M_M}{M_E}}=3.9·10^7\sqrt{\frac{1}{81}}=4.3·10^6\text{ m}."